OpenStudy (anonymous):

if f(x)=2x-3 and g(x)=2x^2+1 find A.f(g(2)) B. g(f(x)) C. does f(x) have and inverse if so find the inverse of f(x)

OpenStudy (anonymous):

f surely has an inverse because it is a line and therefore one to one.

OpenStudy (anonymous):

f says 'multiply by 2 then subtract 3' so f inverse would say to do the opposite things in the opposite order: add 3 and divide by 2 so \[f^{-1}(x)=\frac{x+3}{2}\]

OpenStudy (anonymous):

if this is confusing, rewrite \[f(x)=2x-3\] as \[y = 2x-3\] then switch x and y (because that is what the inverse does) to get \[x=2y-3\] and solve this for x: \[x=2y-3\] \[x+3=2y\] \[\frac{x+3}{2}=y\] and y is your inverse.

OpenStudy (anonymous):

is that the answer for C

OpenStudy (anonymous):

yes. was it clear?

OpenStudy (anonymous):

ok so whats the answers for a and b

OpenStudy (anonymous):

\[f(g(x))=f(2x^2+1)=2(2x^2+1)-3\]

OpenStudy (anonymous):

i guess there is a little algebra to do now: \[2(2x^2+1)-3=4x^2+2-3=4x^2-1\]

OpenStudy (anonymous):

but it was f(g(2))

OpenStudy (anonymous):

is it clear what i did? first write \[f(g(x))\] then replace \[g(x)\] by \[2x^2+1\] and then rewrite f replacing \[x\] by \[2x^2+1\]

OpenStudy (anonymous):

oh well if it is \[f(g(2))\] then since \[f(g(x))=4x^2-1\] then \[f(g(2))=4(2^2)+1=17\]

OpenStudy (anonymous):

typo sorry. \[4(2^2)-1=15\]

OpenStudy (anonymous):

or you could say \[g(2)=2(2^2)+1=9\] and \[f(9)=2\times 9 - 1=15\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[g(f(x))=g(2x-3)=2(2x-3)^2+1\]

OpenStudy (anonymous):

this requires more algebra: \[2(2x-3)^2+1=2(2x-3)(2x-3)+1=2(4x^2-6x+9)+1\] \[=8x^2-12x+18+1=8x^2-12x+19\] if my algebra is correct.

OpenStudy (anonymous):

ok for a how come its 4(2^2)-1

OpenStudy (anonymous):

and not 2(2^2) +1